I am having trouble with my business math homework could someone please help me?
The question is Solve 5/8x + 3/5=8
I did ask one before but I forgot to put all the numbers in sorry.
I would like to thank you in advance for all the help you have giving me so far.

Answer by user_dude2008(1862)

5/8x + 3/5=8


25x + 24=320


25x = 320-24


25x = 296

x=296/25 <-----------------Answer